# Crop Row Switching for Vision-Based Navigation: A Comprehensive Approach for Efficient Crop Field Navigation

### v1

@@ IV-A Crop row exit and headland entry
IV-B U-turn towards next crop row
@@
The U-turn step of the row switching manoeuvre is represented by the state transitions $C\rightarrow D\rightarrow E\rightarrow F$. The angular error for $C\rightarrow D$ transition was calculated based on the angle between $\overrightarrow{AC}$ and $\overrightarrow{DE}$ vectors. The angle between $\overrightarrow{DE}$ and $\overrightarrow{FF_{N}}$ vectors was considered to calculate the angular error for $E\rightarrow F$ transition where $F_{N}$ is a point on the GNSS trajectory which is $N$(=5) points after the GNSS coordinate of state F. Distance error for $D\rightarrow E$ transition was calculated by comparing the $DE$ distance with the inter-row distance between the adjacent crop rows which the robot is being switched, using regressed ground truth lines.
@@
The angular errors in both rotational transitions are evenly distributed with a near zero mean. The absolute median angular errors during rotational transitions were less than 7°. These error margins could be considered as acceptable for this application scenario since such small angular errors wouldn’t incur significant deviations to $\overrightarrow{DE}$ and $\overrightarrow{FG}$ vectors. The robot would stay at the same place without any translational motion in an ideal rotational transition. However, it was evident from some of the recorded trajectories in Figure 5 that these rotations also incur some translational motion within the trajectory. Such motions push the robot towards or away from the direction of next crop row. This would cause the robot to skip the next crop row to be traversed or turn towards the same row it came in, despite achieving the desired $\overrightarrow{DE}$ distance. This unintended translational motion is often caused by the uneven terrain in the headland area which is not detected by the wheel odometry. The headland buffer area of $4^{th}$ crop row had such uneven terrain which lead to failure of both left and right turns originated from that crop row.
IV-C TSM based Re-Entry Validation
@@
The median error for $D\rightarrow E$ transition which led to successful re-entry was always below 30 cm. This indicate that the robot could execute a successful re-entry when it faces the next crop row with a perpendicular offset of 30 cm or below at state $F$. An experiment was setup to validate this hypothesis where the robot was placed facing towards the crop row at different angles within the headland buffer of a given crop row. The TSM algorithm was executed on the robot such that it would detect the crop row in front of it and gradually drive the robot into the crop row in front. The path of the robot was recorded in GNSS coordinates as plotted in Figure 6. All the recorded trajectories could successfully enter the row in front of it since the robot was initiated in the headland buffer and facing towards the general direction of the crop row in front of it.
@@
The re-entry failures in Figure 5 could be explained by the main findings of this experiment. There are two key factors governing the success of a re-entry to the next crop row during $F\rightarrow G$ transition. First requirement is that the robot must be positioned within the headland buffer of the crop row it intends to enter. If the perpendicular offset between the crop row and robot position at state $F$ is beyond 30 cm, a re-entry failure occurs. The second factor is that the robot must be oriented towards the general direction of the crop row it intends to enter. Maximum deviation angle of the robot heading from crop row was 26°in the experiment illustrated in Figure 6. Although the errors in individual state transitions of the row switching manoeuvre are minimal, the overall outcome of all transitions would not lead to success when these two requirements are not met at state $F$.
V Conclusion and Future Work
@@
The proposed row-switching manoeuvre could navigate the robot from one crop row to another without needing RTK-GNSS sensors or multiple cameras while using a single front mounted camera on the robot. Individual steps of the row-switching manoeuvre demonstrated excellent results within the context of each state transition and its functionality. The vision-based $A\rightarrow B$ transition was the transition with highest errors in the proposed manoeuvre. The rotational transitions of the manoeuvre yielded smaller percentage errors relative to the translational transitions. The success rate of the conducted row switching experiment was 55.5% while the re-entry experiment yielded 100% success rate. The row switching manoeuvre is also crop agnostic since it doesn’t rely on plant specific visual features.

### v2

@@ 4.2.1 Crop row exit and headland entry
4.2.2 U-turn towards next crop row
@@
The U-turn step of the row-switching manoeuvre is represented by the state transitions $C\rightarrow D\rightarrow E\rightarrow F$. The angular error for $C\rightarrow D$ transition was calculated based on the angle between $\overrightarrow{AC}$ and $\overrightarrow{DE}$ vectors. The angle between $\overrightarrow{DE}$ and $\overrightarrow{FF_{N}}$ vectors was considered to calculate the angular error for $E\rightarrow F$ transition where $F_{N}$ is a point on the GNSS trajectory which is $N$(=5) points after the GNSS coordinate of state F. Distance error for $D\rightarrow E$ transition was calculated by comparing the $DE$ distance with the inter-row distance between the adjacent crop rows which the robot is being switched, using regressed ground truth lines.
@@
The angular errors in both rotational transitions are evenly distributed with a near-zero mean. The absolute median angular errors during rotational transitions were less than 7. These error margins could be considered acceptable for this application scenario since such small angular errors wouldn’t incur significant deviations to $\overrightarrow{DE}$ and $\overrightarrow{FG}$ vectors. The robot would stay at the same place without any translational motion in an ideal rotational transition. However, it was evident from some of the recorded trajectories in Figure 14 that these rotations also incur some translational motion within the trajectory. Such motions push the robot towards or away from the direction of the next crop row. This would cause the robot to skip the next crop row to be traversed or turn towards the same row it came in, despite achieving the desired $\overrightarrow{DE}$ distance. This unintended translational motion is often caused by the uneven terrain in the headland area which is not detected by the wheel odometry. The headland buffer area of $4^{th}$ crop row had such uneven terrain which led to the failure of both left and right turns originating from that crop row.
4.2.3 TSM-based Re-Entry Validation
@@
The median error for $D\rightarrow E$ transition which led to successful re-entry was always below 30 cm. This indicates that the robot could execute a successful re-entry when it faces the next crop row with a perpendicular offset of 30 cm or below at state $F$. An experiment was set up to validate this hypothesis where the robot was placed facing towards the crop row at different angles within the headland buffer of a given crop row. The TSM algorithm was executed on the robot such that it would detect the crop row in front of it and gradually drive the robot into the crop row in front. The path of the robot was recorded in GNSS coordinates as plotted in Figure 15. All the recorded trajectories could successfully enter the row in front of it since the robot was initiated in the headland buffer and facing towards the general direction of the crop row in front of it.
@@
The re-entry failures in Figure 14 could be explained by the main findings of this experiment. There are two key factors governing the success of re-entry to the next crop row during $F\rightarrow G$ transition. The first requirement is that the robot must be positioned within the headland buffer of the crop row it intends to enter. If the perpendicular offset between the crop row and robot position at state $F$ is beyond 30 cm, a re-entry failure occurs. The second factor is that the robot must be oriented towards the general direction of the crop row it intends to enter. The maximum deviation angle of the robot heading from crop row was 26 in the experiment illustrated in Figure 15. Although the errors in individual state transitions of the row switching manoeuvre are minimal, the overall outcome of all transitions would not lead to success when these two requirements are not met at state $F$.
4.3 Experiment 3: Field Scale Navigation
Figure 16 demonstrates a side-by-side illustration of the crop structure of a real arable field. The tractor initially enters the field and starts drilling around the perimeter as illustrated in Figure 16.b. Then it fills in the center of the field in a lawnmower pattern. The plants drilled around the perimeter are indicated in blue in Figure 16.b while the plants in the centre are marked in green. The demarcation line between non-parallel blue and green plant regions is indicated with orange arrows in Figure 16.a. The crop row switching behaviour proposed in this work is only applicable in the regions of the field indicated in blue. Attempting to execute the crop row switching behaviour in the green region of a real arable field would cause the robot to drive over the crops in the blue region. A field-scale deployment of this proposed infield navigation scheme in a real-world arable field structured for optimal navigation of tractors is limited due to the field structure explained above. Therefore, the field coverage analysis experiment was conducted in a simulated field where the entire field constitutes parallel crop rows with headland space on either side of the ends of the crop rows.
Figure 16: Field structure of a real arable field. a: A satellite image of a real arable field (Copyright: Google Maps), b: The crop structure in a real arable field
4.3.1 initial turning direction detection Analysis
The initial turning direction detection algorithm described in Section 3.6 was tested with the test dataset of the CRDLD dataset. There were 500 images in the test dataset and 10 of them were images of crop rows next to the field edge. The algorithm was able to successfully detect 100% of the images next to the field edge with the correct initial turn direction. However, one false positive was detected during the testing as illustrated in Figure 17. This image was captured in a crop row in the middle of the field. The predicted crop row mask only illustrated the top portions of the crop rows detected on the right side of the central crop row leading the algorithm to predict a false positive edge row.
Figure 17: Field structure of a real arable field. a: A satellite image of a real arable field (Copyright: Google Maps), b: The crop structure in a real arable field.
4.3.2 Field Coverage Analysis
The aim of this experiment is to explore the ability of the proposed scheme to navigate an entire arable field while alternating between row-following and row-switching behaviours. To this end, a simulation environment was set up in the Gazebo simulator with sugar beet plants with realistic textures and the ground texture of the field was transferred from the real environment to the simulation. The parameters pertaining to this simulation environment are listed in Table 5. Plant height and orientation were randomized within the stated variances to mimic realistic variations among the plants in the field as shown in Figure 18.
Table 5: Simulation Parameter for Sugar Beet Field
Property Value Variance
Row Spacing 60 cm 0
Seed Spacing 16 cm 0
Plant Height 6 cm +3 cm
Plant Orientation 0 $\pm$ 145
Row Length 6 m 0
Row Count 20 0
Figure 18: Simulated sugar beet field with the Husky robot at the starting position of the crop row.
This experiment validates initial turning direction detection algorithm introduced in Section 3.6 and the impact of row following and row switching behaviour transitions on field coverage of the proposed infield navigation scheme. The robot will first initiate its infield orientation algorithm to identify the turning direction during its first crop row-switching manoeuvre and continue with row-following behaviour. It will start the row-switching behaviour when it detects its approach towards the end of the row and continues with alternating behaviours until the last crop row. It will stop the navigation scheme when it reaches the last crop row based on the initial turning direction detection algorithm. The robot was positioned at each corner of the simulated sugar beet field similar to its position in Figure 18 and the autonomous navigation scheme was initiated.
Figure 19: Autonomous field scale navigation trajectories for simulated sugar beet field navigation trials. (Each trial starts from each corner of the field indicated by starting points S1-S4).
The autonomous navigation trajectories of the robot during each of the four trials are visualised in Figure 19. The robot has traversed 80 crop rows in total where it failed to traverse 6 crop rows. Therefore the percentage field coverage of all the trials was 92.5%. There were 2 instances in trial 3 in which the robot re-enters back into the same crop row it traversed. Therefore the traversal overlap percentage in the experiment was 2.5%. These errors originated during the $E\rightarrow F$ transition of the crop row switching manoeuvre illustrated in Figure 3. The raw skipping occurs when the robot underturns during $E\rightarrow F$ transition causing the robot to orient itself to the adjacent crop row rather than the intended crop row as seen in row skipping instances of trial 3 of Figure 19. This causes the robot to traverse further than the intended inter-row distance to reach the next crop row to be traversed. The traversal overlap events occur when the robot overturns the expected 90 angle during $E\rightarrow F$ transition, turning back towards the next crop row to be traversed. The row-following algorithm latches the robot towards the already traversed crop row since it is oriented towards the same crop row it already traversed. Both of the described anomalies could be corrected by improving the turning behaviour of the robot turning stage. This could be achieved by employing a predictive controller to avoid overshoots and undershoots, and using sensor fusion for additional positional validation. However, these overshoots and undershoots were not prominent during the turning stages of the real-world crop row-switching experiments conducted in this work. The delay in updating the simulation due to limited computing resources of the system on which the simulator was run could explain the under-performance observed during the simulation. There were two instances where the human driver manually intervened to re-align the robot at the 9th crop row from left during trials 1 and 2.
5 Conclusions and Future Work
A vision-based arable field navigation scheme is presented in this paper. The proposed approach combines alternating robot behaviours for crop row following and crop row switching to achieve field scale navigation. A initial turning direction detection algorithm was proposed to autonomously determine the initial turning direction for a robot starting the field traversal from any corner of the field without pre-assigned directions. The row following behaviour was tested extensively under varying field conditions and growth stages. The autonomous row following behaviour of the robot was tested during a 4.5 km infield navigation experiment which yielded an average of 3.32 cm median cross-track error and 1.24 median heading error. The row following algorithm was challenged due to larger (>1 m) discontinuities in the crop rows where human intervention was needed to re-orient the robot to the crop row.
@@ 5 Conclusions and Future Work
The proposed row-switching manoeuvre could navigate the robot from one crop row to another without needing RTK-GNSS sensors or multiple cameras while using a single front-mounted camera on the robot. Individual steps of the row-switching manoeuvre demonstrated excellent results within the context of each state transition and its functionality. The vision-based $A\rightarrow B$ transition was the transition with the highest errors in the proposed manoeuvre. The rotational transitions of the manoeuvre yielded smaller percentage errors relative to the translational transitions. The success rate of the conducted real world row-switching experiment was 55.5% while the re-entry experiment yielded a 100% success rate. In contrast, the success rate of the simulated row switching experiment (Section 4.3) was recorded at 89.5%.The row-switching manoeuvre is also crop agnostic since it doesn’t rely on plant-specific visual features.
The row following behaviour was combined with the vision-based crop row switching algorithm in a simulated field scale navigation experiment recording a 92.5% field coverage and 2.5% crop row traversal overlap. The row skipping and traversal overlap errors were mainly caused due to the inaccurate rotational movements during the crop row switching manoeuvre. The initial turning direction detection algorithm was tested on the CRDLD dataset with a 100% success rate and re-validated in the simulated deployment of field scale navigation. The difference in success rates of row switching manoeuvre in simulation and the real world could be accounted for by the uneven terrains in the real world and the lack of IMU fused odometry in the robot used in the real-world experiment.
The larger discontinuities are a hard barrier for the current system as the crop row detection and row following commands are generated frame by frame to navigate the robot along the crop row. While complementary filters are implemented to filter out sudden noise in the detected crop row, those filters are not sufficient to prevent the robot from wandering off the traversing crop row path when large discontinuities are encountered. The main problem in overcoming the challenge of larger discontinuities is that the system is unable to distinguish between a large gap in the crop row and the end of the crop row. The information provided by the vision system in both scenarios is largely similar. A software-based solution for such large crop row gaps is to generate a virtual navigation line based on the presence of adjacent crop rows. However, this solution is not very promising in practical settings mainly because large gaps in crop rows are associated with missing crops due to tramlines or foraging wildlife. The likelihood of having lateral crop rows intact in such scenarios is very low. A hardware solution could be suggested with higher confidence to address this problem where the camera could be tilted upwards to capture a larger section of succeeding crop row within the image frame.
@@
The vision-based infield navigation scheme proposed in this thesis was able to achieve 92% field coverage and 2.5% traversal overlap as reported in Chapter 7. However, further development would be necessary in order to reach the ideal 100% field coverage and 0% traversal overlap. While row-following behaviour delivered promising results, crop row-switching algorithms faced significant challenges navigating the headland area mainly using wheel odometry. The errors and challenges encountered during the headland traversal could be mitigated by incorporating additional modalities for localisation during the row-switching manoeuvre. This could be achieved by incorporating visual SLAM algorithms and IMU data during the crop row switching as sensor fusion with the wheel odometry.